What is the net present value of a replacement project whose cash flows are -$856,000; $596,000; $730,000; and $176,000 for years 0 through 3, respectively? The firm has decided to assume that the appropriate cost of capital is 16.9% p.a. (round to the nearest dollar)
Answer #1

Calculate the net present value as follows:

A C D 1 Year Annual cash flows PV factors@16.9% Present value S (856,000.00) 596,000.00 0.855431993156544 S 509,837.47 730,00

A В C D 1 Year Annual cash flows PV factors@ 16.9% Present value -B2*C2 ЕВ3*СЗ =B4*C4 =B5*C5 -SUM(D2:D5) =1/((1+16.9% )^A2) 1